OpenStudy (anonymous):

Medal Reward!!! What is the quotient of three and one over five divided by seven over ten in simplest form? three and two over seven three and seven over fifty four and two over five four and four over seven

OpenStudy (anonymous):

@Juliette32801 this was the hard one.

OpenStudy (calculusxy):

Multiply 16/5x10/7 the product will be your quotient.

OpenStudy (anonymous):

4/7

OpenStudy (anonymous):

@calculusxy gimme a second

OpenStudy (calculusxy):

@Kenx_Khaos I don't know what you mean.

OpenStudy (mathmate):

three and one over five = \( 3\frac{1}{5} \) seven over ten = \( \frac{7}{10}\). So three and one over five divided by seven over ten = \( 3\frac{1}{5} \div \frac{7}{10}\) = \( \frac{16}{5} \times \frac{10}{7}\) =?

OpenStudy (anonymous):

I got 4/7

OpenStudy (anonymous):

I was right! lol

OpenStudy (anonymous):

@calculusxy what I meant by gimme a second I meant hold on

OpenStudy (calculusxy):

Ok.

OpenStudy (anonymous):

lol

OpenStudy (calculusxy):

Sorry as a British, I don't actually understand these things.

OpenStudy (mathmate):

:( 4/7 is not correct, and it's not one of the choices.

OpenStudy (anonymous):

o.o 4 4/7

OpenStudy (mathmate):

That looks a lot better! :)

OpenStudy (anonymous):

I had to add the whole